At the end of 2​ years, P dollars invested at an interest rate r compounded annually increases to an​ amount, A​ dollars, given by the following formula.
A=P(1+r)2

Find the interest rate if ​$128 increased to ​$242 in 2 years.

Answer by stanbon(75887)

Find the interest rate if ​$128 increased to ​$242 in 2 years
--------
A(t) = P(1+(r/n))^(nt)
242 = 128(1+(r/1))^(1*2)
----
(1+r)^2 = 1.8906
1+r = 1.375
rate = 0.375 = 37.5%
------------------
Cheers,
Stan H.